OpenStudy (anonymous):

Simplify Completely: x^2-3x-54/x^2-18x+81 * x^2+12x+36/x+6 a. x+6/(x-9)^2 b.(x+6)^2/x-9 c.(x+6)^2/(x-9)^2 d.x+6/x-9

OpenStudy (cwrw238):

= (x-9)(x+6) * (x+6)(x+6) --------- --------- (x-9)(x-9) (x+6) = (x+6)^2 / (x-9) b

OpenStudy (anonymous):

thank you

OpenStudy (cwrw238):

yw